OpenStudy (anonymous):

how do you slove 4a^4-5a^2+1=0

OpenStudy (amistre64):

solve for a subsituted variable really; claim that x=a^2 and x^2 = a^4 then solve for x... 4x^2 -5x +1 = 0 (x-1/4)(x-4/4) x = 1/4 and x = 1

OpenStudy (amistre64):

since x = a^2 1/4 = a^2 +-sqrt(1/4) = a +- 1/2 = a

myininaya (myininaya):

very good you super hero!

OpenStudy (amistre64):

1 = a^2 1 = a

OpenStudy (amistre64):

lol ..... that aint kyptonite you got in there is it :)

myininaya (myininaya):

it doesnt matter if you arent superman

OpenStudy (amistre64):

they call me ..MISTER PIG!!!....AAAAAAAHHHHHHHHHHHHHH!!!!!!!!!!

OpenStudy (amistre64):

so a = +- 1/2 and or +- 1 and maybe some other things that I forgot to remember..

OpenStudy (anonymous):

thanks!!